Tim and his mother have the same birthday. When Tim is 12, his mother is 48.
How old will they be when his mother's age is 3 times hers? Let this happen in "n" years.

Answer by Boreal(15235)

in n years, his mother will be 48+n and Tim 12+n
(48+n)/(12+n)=3
multiply both sides by (12+n)
48+n=36+3n
12=2n
n=6 years. She will be 54 and he will be 18.